OpenStudy (anonymous):

(1)PT.Which word phrase represents the variable expression? 2n + 4 A. twice a number decreased by 4 B. four more than twice a number C. two more than 4 times a number D. twice the sum of a number and 4

OpenStudy (anonymous):



OpenStudy (anonymous):

Which is an equation for the given word problem? After Elise deposits $15.50 into her bank account, her balance is $325.00. Which equation can be used to find the amount, n, in her account before she made the deposit? A. 15.50 + n = 325.00 B. 15.50 – n = 325.00 C. n + 325.00 = 15.50 D. n – 325.00 = 15.50 Question Resources